If the bottom right corner of an isosceles triangle is 85 degrees what is the vertex?
Answer by stanbon(75887)

If the bottom right corner of an isosceles triangle is 85 degrees what is the vertex?
-----
Draw the picture.
The bottom right AND the bottom left angles are each 85 degrees
----
Note: Sum of the angles of a triangle = 180 degrees.
-----
Equation:
vertex angle = 180-(2*85) = 180-170 = 10 degrees
--------------
Cheers,
Stan H.
----------